### v1

@@ 2.6 Reasoning tasks with Large Language Models
Recent research [41, 40, 42, 43, 44, 45, 46] has investigated the ability of LLMs to function as In-Context Learners (ICLs), meaning they can solve novel and unseen tasks without requiring additional fine-tuning. When provided with a query prompt that includes a context of input-output examples, LLMs can understand the problem and generate appropriate solutions within this framework. LLMs as ICLs have been applied to a wide range of tasks, including planning [41, 46], programming [40, 42], logical problem-solving [43], and symbolic reasoning [44]. In addition to ICL, other paradigms, such as Chain-of-Thought (CoT) [47] and Automatic Chain-of-Thought (ACoT) [23], have been employed to enable LLMs to reason through their responses step-by-step, improving their ability to plan and articulate their thought process explicitly.
Previous research has shown that LLMs can generate semantically meaningful patterns [44, 48], while other work [49] has explored their in-context learning abilities with semantically unrelated labels, where there is no direct relationship between a token and its meaning. Recent studies [41, 50, 45] have investigated using LLMs for creating task-oriented plans
The approach in [45] is the most similar to ours, as it employs a Socratic method [51] that uses video features to predict subsequent actions with an LLM. Socratic Models (SMs) are a framework that utilizes structured dialogue between pre-existing foundation models, each leveraging its unique capabilities based on its training data distribution. However, while PALM integrates an Action Recognition Model (ARM) and a Vision-Language Model (VLM) to generate text prompts for long-term action anticipation, we diverge by using a dual-branch architecture: one branch dedicated to video recognition for action detection and the other utilizing an LLM for action anticipation.
In our mistake detection pipeline, we integrate In-Context Learning (ICL) and Automatic Chain of Thought (ACoT), utilizing an LLM as part of our action anticipation branch. ACoT enables the model to automatically generate intermediate reasoning steps, breaking the problem-solving process into smaller, logical steps. The approach helps the LLM to externalize its reasoning, enhancing its effectiveness in managing complex, sequential tasks.
@@ 3 Methodology
The proposed system leverages a dual-branch framework that integrates the recognition of procedural steps with anticipation modeling, as shown in Fig. 2. In the following sections, we elaborate on the problem formalization (Sec. 3.1), present the branches for step-recognition (Sec. 3.2) and step-anticipation (Sec. 3.3) and finally illustrate the mistake detection procedure (Sec. 3.4).
Figure 2: Our proposed model is based on two main components: The recognition module (orange) processes the input video in an online fashion and predicts actions observed at each timestep; the anticipation module (blue) reasons symbolically via a Large Language Model, utilizing automatic Chain of Thought (ACoT) reasoning to predict the future action based on past action history and a brief context, such as instances of other action sequences. Mistakes are identified when the current action detected by the step recognition method differs from the one forecasted by the step anticipation module.
3.1 Background
Let us consider a finite collection of $N$ procedures $\{p_{i}\}_{i=1}^{N}$, where each procedure encodes a sequence of actions as $p_{i}=\{a_{k}\}_{k=1}^{K_{i}}$. Here, $K_{i}$ varies depending on the specific procedure, and each action $a_{k}$ belongs to the set $a_{k}\in\mathcal{A}=\{a|a\text{ is a possible action}\}$. Additionally, each procedure is represented by a set of videos $\{v_{i}\}_{i=1}^{N}$, which consist of frames $v_{i}=\{f_{\tau}\}_{\tau=1}^{M_{i}}$, with $M_{i}$ indicating the total number of frames in video $i$. Given a specific frame $f_{\tau}$ from video $v_{i}$, our task is two-fold: we (1) identify the action $a_{\tau}$ corresponding to the frame $f_{\tau}$ and (2) predict the action $a_{\tau}$ that will occur at time $\tau$, relying solely on past observations up to time $\tau-1$.
The step recognition task is handled by a module $\rho$, which takes the encoded frames of $v_{i}$ from the start up to frame $\tau$ as input and outputs an action $a^{\rho}_{\tau}$ (Sec. 3.2). Subsequently, we input all the actions $a^{\rho}_{1},...,a^{\rho}_{\tau-1}$ into module $\xi$, which is responsible for the anticipation task, allowing it to predict the next action in the sequence as $a^{\xi}_{\tau}$ (Sec. 3.3).
Finally, we compare $a^{\rho}_{\tau}$ with $a^{\xi}_{\tau}$ designating actions as mistaken when there is a misalignment between the outputs from the two branches (Sec. 3.4). For clarity, the rest of this section will focus on a single procedure $p$ associated with a video $v$.
@@ 3.2 Step Recognition
The step recognition module, referred to as $\rho$, takes a window $W=\{{f_{{\tau}-w},..,f_{\tau}}\}$ with fixed length $w$ of encoded frames from $v$ up to the current frame $\tau$ as input, and is tasked to recognize the last step performed $a^{\rho}_{\tau}$. In our approach, we use MiniRoad [32], the state-of-the-art model for online step detection, which is also computationally efficient in terms of GFlops and parameter count.
In this setup, the model makes predictions for the action $a_{\tau}$ by inspecting the frames ${f_{{\tau}-w},..,f_{\tau}}$.
$\rho({f_{{\tau}-w},..,f_{\tau}})=a^{\rho}_{\tau}$ (1)
This technique often results in redundant predictions, with the model frequently assigning the same action to consecutive frames. To address this issue, we explore methods for aggregating redundant and similar actions in Sec. 6.
During training, the loss for this step recognition module is calculated through a Cross Entropy Loss, comparing the actual action $a_{\tau}$ with the predicted action $a^{\rho}_{\tau}$.
@@ 3.3 Step Anticipation
We employ an LLM as our $\xi$ module for next-step prediction, feeding it with prompts from procedural video transcripts and exploiting its reasoning ability enhanced with the Automatic Chain of Thought (ACoT) [23] technique. Our framework operates without specific training or fine-tuning, leveraging only the reasoning from ACoT (see Fig. 2) and few-shot prompts. These prompts consist of two components: the contextual transcripts $C$ and the current sequence $s_{\tau}$. The former is derived from similar procedures (i.e., different procedures for the same toy) to provide the LLM with information about typical step sequences and their order. The latter, Sequence $s_{\tau}$, includes the current sequence of actions up to the current frame $f_{\tau}$, as they have been detected by our module $\rho$, i.e., the sequence $s_{\tau}$ up to frame $f_{\tau}$ is defined as $s_{\tau}=[a^{\rho}_{1},...,a^{\rho}_{\tau-1}]$, where $a^{\rho}_{i}$ represents the action detected by the module $\rho$ at frame $i$.
@@
The prompting occurs in two stages. First, we employ an ACoT mechanism of $\xi$ as an intermediate step $\phi_{\tau}$. This ACoT process uses $C$, $s_{\tau}$, and an additional prompt to explicitly stimulate reasoning. Then, we use the output of $\phi_{\tau}$ along with $C$ and $s_{\tau}$ to predict the next most probable action. By breaking down the task into smaller logical steps, the LLM can leverage the contextual transcript $C$ and the current action sequence $s_{\tau}$ to generate intermediate reasoning steps, bridging the gap between observed and future actions.
3.4 Mistake Detection
Ultimately, we analyze the outputs of the two modules to identify procedural errors. Specifically, we classify steps as correct when the outputs of both modules match, while we label cases as errors when the outputs differ. That is:
$\begin{cases}a^{\rho}_{\tau}\neq a^{\xi}_{\tau}&\text{MISTAKE}\\ a^{\rho}_{\tau}=a^{\xi}_{\tau}&\text{CORRECT}\\ \end{cases}$ (2)
We compare two evaluation settings: aggregated and frame-level. In both cases, the input data to the recognizer is aggregated; this is done by compressing multiple window frames’ actions in a single one by considering the most frequent action and then merging adjacent actions. While in the aggregated case, we directly compare the anticipator’s output to that aggregated input, in the frame-level case, we have to expand it to the original dimensionality.
4 Benchmark Setup
This section describes the benchmark datasets and evaluation metrics used in our experiments. We introduce Assembly101-O and Epic-tent-O as adaptations of the original datasets [2, 1], detailing the selected labeling criteria for online benchmarking, a novel configuration of training and test splits to accommodate open-set procedural mistakes, and defining our proposed online metrics in Sec. 4.1.
4.0.1 Assembly101-O
Assembly101 [2] is a large-scale video dataset that facilitates procedural video understanding. The dataset contains 362 videos of assembly and disassembly tasks involving 101 types of toy vehicles, recorded from multiple static (8) and egocentric (4) cameras. It includes extensive annotations at various levels of granularity, with over 100K coarse and 1M fine-grained action segments and 18M 3D hand poses. The dataset supports diverse challenges, including action anticipation and segmentation, mistake detection, and 3D pose-based action recognition.
Assembly101 for Online and Open-Set Mistake Detection (Proposed) We introduce a novel split of the dataset [2] tailored explicitly for online, open-set mistake detection. Assembly101-O introduces two main modifications to [2] as we show in Figure 3: a new train/test split and revised procedure lengths. The new split assigns all correct procedures to the training set while reserving videos with mistakes for validation and testing. This adjustment enables models to learn sequences characteristic of correct procedures in a one-class classification (OCC) framework, avoiding bias from specific mistake types during training. By training exclusively on correct sequences, models adhere to the OCC protocol, treating all deviations as mistakes and enabling better balance in the validation and test sets for more comprehensive mistake detection evaluation.
Figure 3: The modification occurs in two steps: starting from Assembly101 (a), first, a new train/test split assigns all correct procedures to the training set, reserving videos with mistakes for validation and testing (b). Second, the lengths of procedures containing mistakes are adjusted by trimming them to the first mistake, preventing the creation of corrupted sequences (c). This setup enables models to learn correct sequences within a one-class classification (OCC) framework, treating any deviation as a mistake and ensuring a balanced validation and test set for effective mistake detection.
The second adjustment limits each procedure’s evaluation to the point where a mistake caused by an incorrect action dependency disrupts the workflow. Per the OCC protocol, models are trained to recognize the correct step sequences necessary for completing procedures. Evaluating beyond the point of procedural compromise creates discrepancies between the training and test set actions, hindering effective step recognition and anticipation. Furthermore, to align with real-world applications, this work focuses on egocentric videos, utilizing only a single egocentric view per procedure from the four available in [2].
4.0.2 Epic-tent-O
Epic-tent is a dataset of egocentric videos capturing the outdoor assembly of a camping tent. It was collected from 24 participants who wore two head-mounted cameras (GoPro and SMI eye tracker) while performing the task, resulting in 5.4 hours of recorded video. The dataset includes annotations for action labels, task errors, participant self-rated uncertainty, and gaze positions, capturing the variability and complexity of the task as participants interacted with non-rigid objects (such as the tent, guylines, instructions, and tent bag) and displayed varying proficiency and confidence levels.
Unlike Assembly101 [2], Epic-tent is designed explicitly for supervised error detection, with each recorded procedure containing at least one mistake, making it incompatible with the split procedure proposed for Assembly101-O. However, Epic-tent provides frame-level confidence scores from each participant, reflecting their self-assessed uncertainty during task performance. Therefore, we define a custom split strategy (detailed in Sec. C of the supplementary materials): videos from higher-confidence participants form the training set. In contrast, participants showing more significant uncertainty (and thus potentially prone to more errors) make up the test set. This strategy is particularly promising for real-world applications where accurately labeling erroneous frames may be challenging or training a mistake detector can start soon after recording without requiring full annotation completion. This partition results in 14 training videos and 15 test videos.
Epic-tent for Online and Open-Set Mistake Detection (Proposed) This section introduces a novel split for the Epic-tent dataset [1] tailored for open-set mistake detection. The dataset includes annotations for nine distinct mistake types, though only four of these categories—“order”, “omit”, “correction”, and “repeat”—are considered procedural mistakes, as they represent deviations that compromise the task’s procedural integrity. Other categories such as “slow”, “search”, “misuse”, “motor”, and “failure” are not procedural errors, as they do not disrupt the task’s progression.
The Epic-tent dataset exclusively comprises egocentric videos recorded with Go-Pro cameras, reinforcing the dataset’s suitability for open-scene, real-world applications. Videos in the test set are trimmed to the frame, marking the first procedural mistake, while videos representing correct procedures remain unaltered.
4.1 Metrics Definition
We work with a sequence of frames, each representing a procedure, where all actions are correct except for the last one, which is a mistake. A mismatch between the prediction of the step recognition and the ones of the step anticipation indicates the presence of a mistake, while no mismatch means the procedure is correct. To assess the performance of our procedural mistake detection model, we follow the evaluation proposed in PREGO: we use True Positives as a measure of the model correctly identifying errors and True Negatives as a measure of accurately labeling steps that are not errors. Thus, we rely on the Precision, Recall, and F1 score metrics to evaluate the performance of our model. These metrics offer valuable insights into the model’s capability to identify and classify mistakes within procedural sequences. Precision quantifies the accuracy when predicting mistakes, minimizing false positives. Recall assesses the model’s capability to retrieve all mistakes, reducing the number of false negatives. Finally, the F1 score is the harmonic mean of precision and recall, and it balances failures due to missing mistakes and reporting false alarms.
To establish a method-agnostic evaluation metric, we focus solely on the model’s ability to classify actions as correct or erroneous. While our approach incorporates an online action detection component, we do not directly assess its general step recognition capabilities—instead, our evaluation centers on the accuracy of distinguishing between correctly and incorrectly performed actions. This decision stems from the limitations of metrics that blend recognizer and anticipator module performances, which can lead to unfair comparisons with approaches employing alternative prediction strategies. Given the expected class imbalance in ego-centric mistake detection tasks and the critical importance of identifying errors, the mistakes are considered as mistakes as the positive class in our evaluation framework.
When evaluating frame-level metrics (see Sec. 7.3), we consider corner cases such as the best, worst, and random scenarios (see Sec. 7.1) for clarity. To ensure a fair comparison, we use balanced accuracy as the evaluation metric, the average of true positive and true negative rates. This is useful to mitigate the impact of dataset imbalance, which would otherwise lead to an unfair comparison.
Balanced $\text{Balanced Accuracy}=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right).$ (3)
5 Large Language Models as Step Anticipators
This section comprehensively analyzes how an action anticipator operates using various techniques and methodologies. We begin by evaluating the performance of several large language models (LLMs) using in-context learning with few-shot examples (Sec. 5.1), which allows us to determine which model performs best in this setting. Once the most effective LLM is identified, we investigate different prompting strategies, starting from a system prompt’s impact, including zero-shot, few-shot, and Automatic-Chain-of-Though prompting (Sec 5.2). In advance, we explore the potential improvements by fine-tuning each prompting method to enhance the model’s performance (Sec. 5.3). Lastly, we perform a speed analysis, comparing different prompting methods (Sec. 5.4).
All experiments use ground truth aggregated labels of the Assembly101-O dataset as the output of the predictor’s branches. This approach ensures that our performance analysis remains objective and untainted by errors from the step recognition branch, allowing for a more accurate evaluation of the step anticipator’s capabilities.
5.1 Model Selection
In this analysis, we evaluate the performance of various LLMs, as shown in Table 2 when using few-shot context. The results demonstrate a clear performance hierarchy among the tested models, with LLAMA 3.1 8B emerging as the superior performer across all metrics. LLAMA 3.1 8B achieved the highest accuracy at 47.40% and the best F1 score of 48.40%, significantly outperforming its counterparts. The performance trend correlates with model size and recency, suggesting that larger, more recently developed models benefit from architectural improvements and more extensive pre-training. This is exemplified by the LLAMA 3.1 8B model, which leverages its 8 billion parameters to capture more complex patterns and better generalize the assembly task. The superior performance of LLAMA 3.1 8B could be attributed to several factors, including potential advancements in its training methodology, the quality and diversity of its training data, and possible optimizations for general task comprehension. An exciting pattern observed across all models is the substantial disparity between Precision and Recall scores. The consistently high Recall scores, peaking at 91.30% for LLAMA 3.1 8B, indicate that these models identify relevant assembly steps. However, the markedly lower Precision scores suggest a tendency towards overprediction, resulting in false positives. This discrepancy highlights the inherent complexity of the assembly prediction task, where models must balance comprehensive step identification with selective precision.
Table 2: Results of different LLMs on Assembly101-O.
Model Accuracy Precision Recall F1
Mistral 7B 35.20 28.60 88.90 45.60
Gemma 9B 31.90 27.20 85.60 43.50
Phi 3 medium 4k instruct 39.60 29.50 88.70 46.20
LLAMA 2 7B 36.3 30.7 94.0 46.3
LLAMA 3 7B 36.40 29.80 87.30 45.60
LLAMA 3.1 8B 47.40 33.80 91.30 48.40
[…]
5.2 Prompt Analysis
System prompt.
The results in Table 3 highlight the positive impact of combining Automatic-Chain-of-Though (ACoT) reasoning with Few-Shot (FS) examples on the performance of LLAMA 3.1 8B across different modalities in the Assembly101-O dataset. In all prompting scenarios, the system prompt is written as:
{myframe}
System: I am going to provide an input sequence that represents a sequence of actions. Your task is to predict the next action of the last sequence based on the patterns observed in the provided input. Limit yourself to only answer with the predicted sequence, and follow the same format given as input.
This prompt consistently offers essential context about the task, improving the model’s performance, particularly in accuracy and precision.
In Zero-Shot (ZS) settings, the instruction prompt is presented as:
{myframe}
@@ 5.2 Prompt Analysis
System: Below is an instruction that describes the task, paired with an input. Write a response that appropriately completes the request.
{myframe}
Toy: ”bulldozer”
Input Sequence:
attach-cabin, attach-body, attach-track, attach-blade
Next Symbol:
attach-figurine
We explore two different input representations: a textual and a numerical one: in the first, we represent an action by its name and natural language, e.g., attach-wheel; in the second one, we use the index associated with the action triplet. For instance, in the textual modality, accuracy improves by 6.8%, and precision increases from 31.80% to 33.80%, reflecting a 6.3% gain. Similarly, in the numerical modality, including the system prompt boosts accuracy by 8.2% (from 40.10% to 43.40%) and precision by 4.6% (from 30.40% to 31.80%).
The improvements can be attributed to the system prompt and FS examples, which provide explicit context and structure, enabling the model to better understand the task’s nuances and generate more accurate and relevant predictions.
Table 3: Results of LLAMA 3.1 8B on Assembly101-O adding a system prompt.
System Prompt Modality Accuracy Precision Recall F1
Textual 47.40 33.80 91.30 48.40
Numerical 43.40 31.80 92.30 46.20
X Textual 44.40 31.80 87.30 46.10
Numerical 40.10 30.40 93.30 46.60
Prompting Methods.
When comparing ZS, FS, and ACoT prompting methods, ACoT emerges as the clear winner (Table 4), particularly when using the pre-trained model (Base), across both textual and numerical tasks. In the ACoT approach, automatic ACoT reasoning is employed, where the LLM is queried twice: first, to generate an internal reasoning step for the answer, and then to use this reasoning, combined with a few examples (as in FS), to derive the final answer. The first query is:
{myframe}
@@
    For the action ”attach the cabin,” consider the immediate effect and long-term consequences.

### v2

@@ 2.5 Reasoning Tasks with Large Language Models
Previous research has shown that LLMs can generate semantically meaningful patterns [46, 42], while other work [47] has explored their in-context learning abilities with semantically unrelated labels, where there is no direct relationship between a token and its meaning. More recent work  [48, 45, 49, 50, 46, 51, 52] has further investigated the ability of LLMs to function as In-Context Learners (ICLs), meaning they can solve novel and unseen tasks relying only on a structured and informative prompt, without requiring additional fine-tuning. When provided with a query prompt that includes a context of input-output examples, LLMs can understand the problem and generate appropriate solutions within this framework. LLMs as ICLs have been applied to a wide range of tasks, including planning [48, 52], programming [45, 49], logical problem-solving [50] and symbolic reasoning [46]. In addition to ICL, other paradigms, such as Chain-of-Thought (CoT) [53] and Automatic Chain-of-Thought (ACoT) [28], have been employed to enable LLMs to reason through their responses step-by-step, improving their ability to plan and articulate their thought process explicitly.
Palm [51] is the most similar to our anticipation module, as it employs a Socratic method [54] that uses video features to predict subsequent actions with an LLM. Socratic Models are a framework that utilizes structured dialogue between pre-existing foundation models, each leveraging its unique capabilities based on its training data distribution. However, while PALM  [55] integrates an Action Recognition Model and a Vision-Language Model to generate text prompts for long-term action anticipation, we diverge by using a dual-branch architecture: one branch tasked to recognize the current action and the other leveraging an LLM for action anticipation focusing on next step prediction. This schema offer a natural pipeline for error detection by comparing the predictions generated by the two branches
In our mistake detection pipeline, we integrate In-Context Learning and Automatic Chain of Thought, utilizing an LLM as part of our action anticipation branch. ACoT enables the model to automatically generate intermediate reasoning steps, breaking the problem-solving process into smaller, logical steps. The approach helps the LLM to externalize its reasoning, enhancing its effectiveness in managing complex, sequential tasks.
@@ 3 Methodology
The proposed system leverages a dual-branch framework that integrates the recognition of procedural steps with anticipation modeling, as shown in Fig. 2. In the following sections, we elaborate on the problem formalization (Sec. 3.1), present the branches for Step Recognition (Sec. 3.2) and Step Anticipation (Sec. 3.3) and finally illustrate the Mistake Detection procedure (Sec. 3.4).
Figure 2: Our proposed model is based on two main components. The recognition module (orange) processes the input video in an online fashion and predicts actions observed at each timestep represented as a square. This prediction is combined with previous ones to minimize the noise generated by per-frame predictions. The last element of the aggregated output corresponds to the current step predicted by the recognition module, while the sequence without duplicates is provided to the anticipation module. The anticipation module (blue) reasons symbolically via a Large Language Model, utilizing automatic Chain of Thought (ACoT) reasoning to predict the future action based on past action history and a brief context such as instances of other step sequences. Mistakes are identified when the current action detected by the step recognition method differs from the one forecasted by the step anticipation module.
3.1 Preliminaries
Let us consider a finite collection of $N$ procedures $\{p_{i}\}_{i=1}^{N}$, where each procedure encodes a sequence of steps as $p_{i}=\{s_{k}\}_{k=1}^{K_{i}}$ that represents the transcript of $p_{i}$. Here, $K_{i}$ varies depending on the specific procedure, and each step $s_{k}$ belongs to the set of all possible steps $\mathcal{S}$ in the dataset. Furthermore, each procedure is represented by a set of videos $\{v_{i}\}_{i=1}^{N}$, which consist of frames $v_{i}=\{f_{\tau}\}_{\tau=1}^{M_{i}}$, with $M_{i}$ indicating the total number of frames in video $i$. Given a specific frame $f_{\tau}$ from video $v_{i}$, we: (1) classify the step $s_{\tau}$ corresponding to the frame $f_{\tau}$ (i.e., step recognition) and (2) predict the step $s_{\tau}$ that occurring at time $\tau$, relying solely on the previously recognized steps up to time $\tau-1$ (i.e., step anticipation).
Step recognition is handled by a module $\rho$, which takes the encoded frames of $v_{i}$ up to the $\tau$-th frame as input and outputs the recognized step $s^{\rho}_{\tau}$ (Sec. 3.2). Subsequently, all recognized steps $s^{\rho}_{1},...,s^{\rho}_{\tau-1}$, once aggregated, are fed into the module $\xi$, which handles the anticipation task by predicting the next step in the procedure, referred to as $s^{\xi}_{\tau}$(Sec. 3.3).
Finally, we compare $s^{\rho}_{\tau}$ with $s^{\xi}_{\tau}$ and deem the step as mistaken if there is a misalignment between the outputs from the two branches (Sec. 3.4). For clarity, the rest of this section will focus on a single procedure $p$ associated with a video $v$.
@@ 3.2 Step Recognition
The step recognition module, denoted $\rho$, takes a window $W=\{{f_{{\tau}-w},..,f_{\tau}}\}$ with a fixed length $w$ of frames from $v$ to the current frame $\tau$ as input and is tasked with recognizing the last step performed $s^{\rho}_{\tau}$. We leverage MiniRoad [37] as $\rho$, which currently represents the state-of-the-art and is also computationally efficient in terms of GFlops and parameter count.
In this setup, the model predicts step $s_{\tau}$ by analyzing the frames ${f_{\tau-w},..,f_{\tau}}$.
$\rho(f_{\tau-w:\tau})=s^{\rho}_{\tau}$ (1)
During training, the loss for $\rho$ is calculated using a cross-entropy loss, comparing the actual step $s_{\tau}$ with the predicted one $s^{\rho}_{\tau}$ as detailed in [37].
Step aggregation
Despite being a state-of-the-art model, MiniRoad [37] performs suboptimally on procedural datasets, primarily due to two key factors: the shift in action distribution between the training and testing phases, and the inherent class imbalance within the dataset. The distribution shift occurs because the actions seen during training often differ from those encountered during testing, leading to poorer generalization. Additionally, Assembly101 contains a significant imbalance in class frequencies (see Figure 3 for Assembly101-O’s split), where common actions dominate the dataset while less frequent yet crucial procedural actions are underrepresented.
Figure 3: Distribution of the 20 most frequent classes in Assembly101-O
To address this, we propose a frame aggregation strategy that efficiently converts per-frame predictions into coherent action sequences, aiming to reduce noise and stabilize predictions. This results in a cleaner input for the anticipator module, dubbed $AG$ in Fig. 2.
In our experiments, we asses three strategies to perform aggregation. One approach (Non-Overlapping Mode Aggregation, NOMA) uses nonoverlapping windows, while the other two employ overlapping windows. These two overlapping methods differ based on which frame is selected for imputing the aggregation result: the last frame (Overlapping Mode Aggregation, OMA) or the central frame (Overlapping Central Mode Aggregation, OCMA) (See Figure 4). All approaches are designed to reduce noise and eliminate redundant consecutive predictions, ensuring that the resulting sequence aligns closely with the actual procedure measured by the Levenshtein Similarity. In section 5.2, we compare and discuss in details the aggregation strategies and find NOMA to be the most effective, which we therefore select.
Figure 4: The three aggregation strategies operate differently: NOMA employs a non-overlapping window and replaces the entire window content with its mode. In contrast, both OMA and OCMA use overlapping windows with a sliding step of 1. However, OMA replaces the last frame of the window with its mode, whereas OCMA replaces the central frame.
@@ 3.3 Step Anticipation
We employ an LLM as our $\xi$ module for next-step prediction, using as input transcripts from procedural video. Our step anticipation operates without domain specific training or fine-tuning, leveraging only the reasoning from Automatic Chain of Thought (ACoT) [28] (see Fig. 2) and few-shot prompting method. These prompts consist of three components: the system prompt, the ACoT prompt, which enables reasoning, and the contextual transcripts $C$ together with the current aggregated sequence $AG$. $C$ is derived from similar transcripts (i.e., different videos for the object we are considering) to provide the LLM with information about typical step sequences and their order (see following examples). Sequence $S^{\rho}_{\tau}$, includes the current sequence of actions up to the current frame $f_{\tau-1}$, as they have been detected by our module $\rho$, i.e., $S^{\rho}_{\tau}=[s^{\rho}_{1},...,s^{\rho}_{\tau-1}]$, where $s^{\rho}_{i}$ represents the action detected by the module $\rho$ at frame $i$. $S^{\rho}_{\tau}$ then undergoes our aggregation strategy, which results in $AG$.
@@
The prompting occurs in two stages. First, we employ an ACoT mechanism of $\xi$ as a first step. This ACoT process uses context $C$, aggregation sequence $AG$, and an additional prompt to explicitly stimulate reasoning (See AcoT Prompt in the following scheme). Then, we use the output of the intermediate step along with $C$ and $AG$ to predict the next most probable action. By breaking down the task into smaller logical steps, the LLM can leverage the contextual transcript $C$ and the current action sequence $AG$ to generate intermediate reasoning steps, bridging the gap between observed and future actions $s^{\xi}_{\tau}$.
{systemframe}
System Prompt:
@@
Below is an instruction that describes the task, paired with an input. Write a response that appropriately completes the request.
{acotframe}
ACoT Prompt:
@@
    For the action “attach the cabin”, consider the immediate effect and long-term consequences.